### v1

@@ Abstract
Reservoir is an efficient network for time series processing. It is well known that network structure is one of the determinants of its performance. However, the topology structure of reservoirs, as well as their performance, is hard to analyzed, due to the lack of suitable mathematical tools. In this paper, we study the topology structure of reservoirs using persistent GLMY homology theory, and develop a method to improve its performance. Specifically, it is found that the reservoir performance is closely related to the one-dimensional GLMY homology groups. Then, we develop a reservoir structure optimization method by modifying the minimal representative cycles of one-dimensional GLMY homology groups. Finally, by experiments, it is validated that the performance of reservoirs is jointly influenced by the reservoir structure and the periodicity of the dataset.
@@ 1 Introduction
Neural networks, inspired by the biological brain, have become a cornerstone in artificial intelligence and machine learning, enabling solutions to complex tasks in image recognition, natural language processing, and predictive analytics [1, 2]. Their structure—including layer count, neuron connectivity, and activation function type—directly determines functional capabilities, with early studies establishing foundational concepts like weight roles and nonlinearity from activation functions [3]. Moreover, a recent study has proposed the weight-agnostic neural networks [4], which can finish learning tasks without weight training but only with structure optimization. This shows that the structure of neural networks plays an essential role. However, as real-world problem complexity grows, a deeper understanding of task-optimized neural network structure design is needed.
@@
The continuous improvement and optimization of neural network structures hold profound and far-reaching significance in advancing the practical application and theoretical development of artificial intelligence. From a performance perspective, a well-optimized structure can significantly enhance the network’s ability to extract features and learn patterns, enabling it to achieve higher accuracy and robustness in handling complex tasks. In terms of resource efficiency, optimizing the number of layers, neuron density, and connectivity patterns can effectively reduce the model’s parameter scale and computational complexity. Additionally, structural optimization plays a crucial role in enhancing the adaptability and interpretability of neural networks. By adjusting the structure to match the characteristics of specific tasks, the network’s generalization ability across different domains can be improved. Meanwhile, rational structural design can reduce the “black-box” nature of neural networks—for instance, through modular structural design, researchers can more clearly trace the flow of information and the contribution of each module to the final output, laying the foundation for interpreting the internal working mechanism of the network [5]. Moreover, recent studies indicate that the topology of a network’s structure significantly impacts its performance, suggesting that understanding and designing network architectures from a topological perspective is a crucial and effective approach [6].
@@
Reservoir Computing (RC), as a distinctive paradigm within the machine learning landscape, has rapidly gained traction and widespread adoption owing to its inherent advantages that address key challenges in traditional neural network deployment. Specifically, RC exhibits a compact model architecture, which minimizes the demand for storage resources; enables rapid training processes by focusing optimization only on the readout layer rather than the entire network, significantly reducing computational overhead; delivers high prediction accuracy across various time-series and nonlinear tasks; and maintains excellent generalization performance when faced with unseen or noisy data—these attributes collectively make RC a preferred choice for applications ranging from real-time signal processing to complex system prediction [7, 8]. Nevertheless, a critical limitation of RC lies in its “black-box” nature, primarily stemming from the lack of clear interpretability regarding the internal dynamics of the reservoir. The reservoir, typically composed of a large number of randomly connected neurons, operates through complex nonlinear interactions, making it difficult to establish a direct link between its structural characteristics (e.g., connectivity density, neuron type, and topological arrangement) and its prediction performance. Consequently, investigating how the internal structure of the reservoir influences its prediction ability has become an increasingly attractive and essential research direction, as resolving this issue could not only unlock the potential to further optimize RC performance but also enhance its reliability and applicability in critical domains where interpretability is essential [9].
@@
In practice, a reservoir is usually regarded as a digraph. In previous studies, the lack of mathematical tools to study the topology of digraphs made it difficult to topologically optimize the structure of the reservoir from an algebraic topological perspective. Now with the rapid development of GLMY homology theory, there are new effective tools for topology optimization of digraphs. The GLMY homology, also known as path homology and initially proposed by [10], represents a significant development in computational topology. As complex systems—ranging from social networks and neural connections to transportation and biochemical pathways—are increasingly modeled using directed graphs (digraphs), traditional homology theories, which are designed for undirected topological spaces, prove inadequate. They fail to capture crucial directional information, such as information flow, influence propagation, or causal relationships. GLMY homology overcomes this fundamental limitation by constructing a chain complex specifically based on directed paths within digraphs. Its distinct sensitivity to edge directions allows it to capture the flow and hierarchical organization of these paths, revealing hidden structural features and topological properties that are otherwise invisible to standard methods. In real-world digraphs, edge orientation often conveys critical information, such as influence direction in a social network or material flow in a manufacturing process. The directional awareness of GLMY homology provides a more accurate and nuanced representation of the digraph’s underlying topology. Persistent GLMY homology [11], a further development of GLMY homology, enhances its analytical capabilities by tracking how directed path structures evolve as parameters, such as edge weights, change. Through persistent GLMY homology, researchers can effectively capture structural features encoded in both edge directions and weights, making it a powerful tool for analyzing dynamic and weighted complex systems. This framework provides a new lens through which to analyze the structural properties of complex networks, particularly those where directionality is a defining feature.
@@
Therefore, in this study, since a reservoir can be represented as a digraph, we present a novel method for reservoir optimization using GLMY homology. Firstly, we show that adding rings  [^1] While some literature uses the term “cycles”, we adopt “rings” to avoid conceptual conflicts with homology theory. [12, 13] in the reservoir digraph can improve the orthogonality of the reservoir adjacent matrix, hence can enhance the reservoir’s predictive capacity. This serves as the fundamental motivation for our approach. Our strategy focuses on strategically modifying the edge directions within the digraph to deliberately increase the number of rings, thereby optimizing the reservoir’s performance. Moreover, since these rings are a special type of representative cycle of one-dimensional GLMY homology, we can achieve our target by considering a subset of representative cycle of 1-d GLMY homology, i.e., the subset consists of all rings. We also validated by experiments that the prediction ability of a reservoir is related to both its structure and the topological features of the dataset.
@@
  • Adding rings inside the reservoir digraph can improve the reservoir adjacent matrix’s orthogonality, which is equivalent to increase a certain type of representative cycle of GLMY homology.
  • We optimize reservoirs’ structure by expanding a certain subset of representative cycle (rings) of 1-d GLMY homology.
@@ 2 Related Work
In this section, we provide some related work about optimizing reservoir structure, and the GLMY homology and some of its applications.
@@ 2.1 Reservoir computing and reservoir optimization
Reservoir computing (RC) has emerged as a powerful and widely adopted machine learning paradigm over the past few decades, gaining considerable attention for its effectiveness in handling complex time-dependent data. Its application has extended across various domains, including time series prediction, nonlinear system identification [14], and speech recognition [15]. The core strength of RC lies in its architectural simplicity: a fixed, randomly generated recurrent neural network (the reservoir) projects the input into a high-dimensional feature space, while only a simple linear readout layer is trained. This separation of the nonlinear dynamics of the reservoir from the linear readout training process is a key reason for its exceptional computational efficiency and rapid training speed. This efficiency has led to remarkable results, such as those demonstrated by Jaeger et al. [16], who applied RC to chaotic time series prediction and achieved computational accuracy two orders of magnitude higher than traditional approaches. Beyond its core architecture, variations have further expanded its capabilities. For instance, Shougat et al. [17] employed Hopf oscillators as an intermediate reservoir, successfully demonstrating robustness to noise and preservation of predictive accuracy even with reduced training costs. In the financial sector, Wang et al. [18] utilized an RC network for forecasting stock market prices, where comparative experiments against several classical models—including long short-term memory (LSTM) networks—showed the RC model consistently achieving lower mean-squared error (MSE) and mean absolute error (MAE), alongside higher $R^{2}$ scores. This work also provided insights into how distinct reservoir initialization strategies influence predictive accuracy, further highlighting the critical role of network dynamics in financial forecasting.
@@
A reservoir is usually represented as a digraph (called the reservoir digraph). In most implementations, reservoir parameters and topological configurations are initialized stochastically. However, such stochastic initialization frequently produces unstable dynamical behavior, necessitating extensive empirical trial-and-error to obtain a well-conditioned network. Therefore, beyond investigating the deployment of reservoirs in prediction tasks, the design of high-quality reservoir architectures has emerged as both a major challenge and a focal point of contemporary research. Gallicchio et al. [12] assessed the impact of constrained reservoir topologies—namely permutation, ring, and chain structures—within Deep Echo State Networks (DeepESNs) on time series forecasting performance. Through a systematic comparison against conventional sparsely random reservoirs and shallow ESNs, demonstrated that embedding structurally sparse topologies in deep reservoir computing constitutes a straightforward yet effective strategy for enhancing temporal modeling accuracy. Cui et al. [19] investigated the interplay between reservoir topology and predictive performance by constructing reservoirs with small-world, scale-free, and hybrid small-world/scale-free structures. Experimental results demonstrate that manually designed reservoir networks deliver superior predictive accuracy and possess a broader spectral radius, while their short-term memory capacity remains comparable to that of traditional models. Moreover, Roeschies et al. [20] proposed an evolutionary algorithm that optimizes reservoir-network architectures for specific problem classes, resulting in substantial gains in predictive performance. Sun et al. [21] introduced a deterministic echo state network architecture featuring a loop reservoir with adjacent feedback connections. This simplified design builds upon the basic loop reservoir structure by incorporating regular adjacent feedback, requiring optimization of only a single parameter. The approach significantly reduces the complexity of ESN implementation while maintaining computational capabilities. These results, together with previous studies, confirm that the structure of a reservoir is pivotal in determining its functionality and overall effectiveness.
@@ 2.2 GLMY homology and its applications
GLMY homology represents a significant extension of classical homology theory, specifically developed to analyze directed or asymmetric interactions in digraphs [10] and hypergraphs [22]. This framework has spawned several theoretical developments, such as homotopy theory for digraphs [23], discrete Morse theory on digraphs [24], and path homology theory for multigraphs and quivers [25]. Notable computational advances include the persistent path homology for networks proposed by Chowdhury and Mémoli [11], which offers a novel approach to digraph analysis in topological data analysis, and an efficient algorithm for computing one-dimensional persistent path homology of digraph filtrations by Dey et al. [26]. In contrast to traditional simplicial complexes, GLMY homology offers a unique set of dimensional fingerprints. These fingerprints are instrumental in unearthing the intrinsic structural information hidden within networks, providing a more detailed view of network architecture. This theoretical framework has demonstrated its practical utility across a diverse range of applications, such as modeling intricate biological systems [27], analyzing material properties and structures [28], and understanding complex networks [29, 30]. Persistent GLMY homology represents a further advancement in this area. By tracking how topological features change and evolve across various filtration processes, it allows us to identify and analyze topological patterns in networks [11]. This capability significantly enhances our understanding of network behavior and evolution.
@@ 3 Preliminary
A digraph is an ordered pair $G=(V,E)$, where $V$ denotes the vertex set and $E\subseteq V\times V$ represents the directed edges. A digraph without loops or multiple edges is called a simple digraph. The GLMY homology (or path homology) provides a topological framework for analyzing simple digraphs [10]. For a finite nonempty vertex set $V$, an elementary $n$-path is defined as a sequence $e_{i_{0}\ldots i_{n}}$ of $n+1$ vertices, where $n\geq 0$. By convention, the set of elementary (1)(-1)-paths is empty. The boundary operator acts on elementary $n$-paths as:
@@
Figure 2: (a) Example of an one-dimensional persistence diagram. (b) An illustration of the confidence set (red area) of (a).
Persistent homology is a tool from computational topology that captures the topological features of space at various scales [31]. It provides a way to quantify the persistence of topological features, capturing the birth and death of topological features in the considered space (such as point clouds, graphs, and digraphs) by constructing a filtration. The persistence diagram (PD) serves as a powerful visualization tool for persistent homology, where each point $(b_{i},d_{i})$ corresponds to a persistence pair representing the birth ($b_{i}$) and death ($d_{i}$) times of a homology generator (Fig. 2). Notice that all points in a PD are located above or on the diagonal. And the value $|d_{i}-b_{i}|$ is called the persistence of the point $(b_{i},d_{i})$, representing the significance of the corresponding topological features. The greater the persistence of a point in the persistence diagram, the more significant the corresponding topological feature is.
A statistical approach for identifying significant points in a persistence diagram is the confidence set [32], which categorizes these points into two classes based on their persistence. As shown in Fig. 2 (b), the confidence set draws two areas on the persistence diagram: the red area and the white area, which contains points with less persistence and greater persistence, respectively. Points in the class that show greater persistence are called significant points, while the other points are called noise points.
The tool of persistence diagram is first employed to illustrate the persistent homology computed from some point clouds, and this framework can also naturally extend to persistent GLMY homology, where we denote the one-dimensional persistence diagram of $\mathbb{H}_{1}$ as $Dgm_{1}(G)$.

### v2

@@ Abstract
Reservoirs are efficient networks for time-series processing. It is well known that the network structure is one of the determinants of their performance. However, the topological structure of reservoirs, as well as their performance, is hard to analyze due to the lack of suitable mathematical tools. In this paper, we study the topological structure of reservoirs using persistent GLMY homology theory and develop a method to improve their performance. Specifically, we find that reservoir performance is correlated with the one-dimensional GLMY homology groups. Then, we develop a reservoir structure optimization method by modifying the minimal representative cycles of one-dimensional GLMY homology groups. Finally, through experiments, we validate that the performance of reservoirs is jointly influenced by the reservoir structure and the periodicity of the dataset.
@@ 1 Introduction
Neural networks, inspired by biological brains, have become a cornerstone in artificial intelligence and machine learning, enabling solutions to complex tasks in image recognition, natural language processing, and predictive analytics [lecun2015deep, goodfellow2016deep]. Their structure—including layer count, neuron connectivity, and activation function type—directly determines functional capabilities, with early studies establishing foundational concepts like weight roles and nonlinearity from activation functions [hornik1989multilayer]. Moreover, a recent study has proposed weight-agnostic neural networks [gaier2019weight], which can perform learning tasks without weight training, relying only on structural optimization. This shows that the structure of neural networks plays an essential role. However, as real-world problem complexity grows, a deeper understanding of task-optimized neural network structure design is needed.
@@
The continuous improvement and optimization of neural network structures hold profound and far-reaching significance in advancing the practical application and theoretical development of artificial intelligence. From a performance perspective, a well-optimized structure can significantly enhance the network’s ability to extract features and learn patterns, enabling it to achieve higher accuracy and robustness in handling complex tasks. In terms of resource efficiency, optimizing the number of layers, neuron density, and connectivity patterns can effectively reduce the model’s parameter scale and computational complexity. Additionally, structural optimization plays a crucial role in enhancing the adaptability and interpretability of neural networks. By adjusting the structure to match the characteristics of specific tasks, the network’s generalization ability across different domains can be improved. Meanwhile, rational structural design can reduce the “black-box” nature of neural networks—for instance, through modular structural design, researchers can more clearly trace the flow of information and the contribution of each module to the final output, laying the foundation for interpreting the internal working mechanism of the network [montavon2018methods]. Moreover, recent studies indicate that the topology of a network’s structure significantly impacts its performance, suggesting that understanding and designing network architectures from a topological perspective is a crucial and effective approach [zhang2024comprehensive].
@@
Reservoir Computing (RC), as a distinctive paradigm within the machine learning landscape, has rapidly gained traction and widespread adoption owing to its inherent advantages that address key challenges in traditional neural network deployment. Specifically, RC exhibits a compact model architecture, which minimizes the demand for storage resources; enables rapid training processes by focusing optimization only on the readout layer rather than the entire network, significantly reducing computational overhead; delivers high prediction accuracy across various time-series and nonlinear tasks; and maintains excellent generalization performance when faced with unseen or noisy data—these attributes collectively make RC a preferred choice for applications ranging from real-time signal processing to complex system prediction [lukovsevivcius2009reservoir, jaeger2001echo]. Nevertheless, a critical limitation of RC lies in its “black-box” nature, primarily stemming from the lack of clear interpretability regarding the internal dynamics of the reservoir. The reservoir, typically composed of a large number of randomly connected neurons, operates through complex nonlinear interactions, making it difficult to establish a direct link between its structural characteristics (e.g., connectivity density, neuron type, and topological arrangement) and its prediction performance. Consequently, investigating how the internal structure of the reservoir influences its prediction ability has become an increasingly attractive and essential research direction, as resolving this issue could not only unlock the potential to further optimize RC performance but also enhance its reliability and applicability in critical domains where interpretability is essential [pathak2018model].
@@
In practice, a reservoir is usually regarded as a digraph. In previous studies, the lack of mathematical tools to study the topology of digraphs made it difficult to topologically optimize the structure of the reservoir from an algebraic topological perspective. Now with the rapid development of GLMY homology theory, there are new effective tools for topology optimization of digraphs. The GLMY homology, also known as path homology and initially proposed by [grigor2012homologies], represents a significant development in computational topology. As complex systems—ranging from social networks and neural connections to transportation and biochemical pathways—are increasingly modeled using directed graphs (digraphs), traditional homology theories, which are designed for undirected topological spaces, prove inadequate. They fail to capture crucial directional information, such as information flow, influence propagation, or causal relationships. GLMY homology overcomes this fundamental limitation by constructing a chain complex specifically based on directed paths within digraphs. Its distinct sensitivity to edge directions allows it to capture the flow and hierarchical organization of these paths, revealing hidden structural features and topological properties that are otherwise invisible to standard methods. In real-world digraphs, edge orientation often conveys critical information, such as influence direction in a social network or material flow in a manufacturing process. The directional awareness of GLMY homology provides a more accurate and nuanced representation of the digraph’s underlying topology. Persistent GLMY homology [chowdhury2018persistent], a further development of GLMY homology, enhances its analytical capabilities by tracking how directed path structures evolve as parameters, such as edge weights, change. Through persistent GLMY homology, researchers can effectively capture structural features encoded in both edge directions and weights, making it a powerful tool for analyzing dynamic and weighted complex systems. This framework provides a new lens through which to analyze the structural properties of complex networks, particularly those where directionality is a defining feature.
@@
Therefore, in this study, since a reservoir can be represented as a digraph, we present a novel method for reservoir optimization using GLMY homology. Firstly, we show that adding rings  [^1] While some literature uses the term “cycles”, we adopt “rings” to avoid conceptual conflicts with homology theory. in the reservoir digraph can improve the orthogonality of the reservoir adjacency matrix, hence enhancing the reservoir’s predictive capacity. Mathematically, a directed ring structure with uniform weights corresponds to a (scaled) permutation adjacent matrix. Permutation matrices are a fundamental subclass of orthogonal matrices. As highlighted in [gallicchio2019reservoir], structured reservoirs based on permutation matrices provide a particularly advantageous architectural setting by ensuring stable signal propagation and mitigating the vanishing gradient problem through their near-orthogonal nature. Moreover, Rodan and Tino [rodan2010minimum] introduced the Simple Cycle Reservoir (SCR), proving that a deterministic ring topology is sufficient to capture complex temporal features and can reach near-optimal memory capacity. And theoretical studies in [stockdill2016restricted] also confirm that rings are essential for replicating the memory-leaking and state-holding properties required for temporal processing. By increasing the number of rings, one can effectively increase the number of paths for information to recirculate, which provably maintains the “fading memory” property over longer time horizons compared to purely random connections. Hence, these studies serve as the fundamental motivation for our approach. Our strategy focuses on strategically modifying the edge directions within the digraph to deliberately increase the number of rings, thereby optimizing the reservoir’s performance. Moreover, since these rings are a special type of representative cycle of one-dimensional GLMY homology, we can achieve our target by considering a subset of representative cycles of 1D GLMY homology, i.e., the subset consists of all rings. In other words, our optimization strategy, which focuses on increasing the number of rings, is theoretically motivated by the advantages of cyclic topologies. According to the SCR theory [rodan2010minimum], rings provide the necessary feedback loops for stable fading memory. By leveraging GLMY homology, we provide a systematic way to enhance these critical cyclic structures in an initially random reservoir, effectively moving its topology toward a more organized, high-memory-capacity state. We also validate through experiments that the prediction ability of a reservoir is related to both its structure and the topological features of the dataset.
@@
  • Adding rings inside the reservoir digraph can improve the orthogonality of the reservoir adjacency matrix, which is equivalent to increasing a certain type of representative cycle of GLMY homology.
  • We optimize the structure of reservoirs by expanding a certain subset of representative cycles (rings) of 1D GLMY homology.
@@ 2 Related Work
In this section, we provide some related work about optimizing reservoir structure, the GLMY homology, and some of its applications.
@@ 2.1 Reservoir computing and reservoir optimization
Reservoir computing (RC) has emerged as a powerful and widely adopted machine learning paradigm over the past few decades, gaining considerable attention for its effectiveness in handling complex time-dependent data. Its application has extended across various domains, including time series prediction, nonlinear system identification [yao2019prediction], and speech recognition [skowronski2007noise]. The core strength of RC lies in its architectural simplicity: a fixed, randomly generated recurrent neural network (the reservoir) projects the input into a high-dimensional feature space, while only a simple linear readout layer is trained. This separation of the nonlinear dynamics of the reservoir from the linear readout training process is a key reason for its exceptional computational efficiency and rapid training speed. This efficiency has led to remarkable results, such as those demonstrated by Jaeger et al. [jaeger2004harnessing], who applied RC to chaotic time series prediction and achieved computational accuracy two orders of magnitude higher than traditional approaches. Beyond its core architecture, variations have further expanded its capabilities. For instance, Shougat et al. [shougat2021hopf] employed Hopf oscillators as an intermediate reservoir, successfully demonstrating robustness to noise and preservation of predictive accuracy even with reduced training costs. In the financial sector, Wang et al. [wang2021stock] utilized an RC network for forecasting stock market prices, where comparative experiments against several classical models—including long short-term memory (LSTM) networks—showed the RC model consistently achieving lower mean-squared error (MSE) and mean absolute error (MAE), alongside higher $R^{2}$ scores. This work also provided insights into how distinct reservoir initialization strategies influence predictive accuracy, further highlighting the critical role of network dynamics in financial forecasting.
@@
A reservoir is usually represented as a digraph (called the reservoir digraph). In most implementations, reservoir parameters and topological configurations are initialized stochastically. However, such stochastic initialization frequently produces unstable dynamical behavior, necessitating extensive empirical trial‑and‑error to obtain a well‑conditioned network. Therefore, beyond investigating the deployment of reservoirs in prediction tasks, the design of high‑quality reservoir architectures has emerged as both a major challenge and a focal point of contemporary research. Gallicchio et al. [gallicchio2019reservoir] assessed the impact of constrained reservoir topologies—namely permutation, ring, and chain structures—within Deep Echo State Networks (DeepESNs) on time series forecasting performance. Through a systematic comparison against conventional sparsely random reservoirs and shallow ESNs, demonstrated that embedding structurally sparse topologies in deep reservoir computing constitutes a straightforward yet effective strategy for enhancing temporal modeling accuracy. Cui et al. [cui2012architecture] investigated the interplay between reservoir topology and predictive performance by constructing reservoirs with small‑world, scale‑free, and hybrid small‑world/scale‑free structures. Experimental results demonstrate that manually designed reservoir networks deliver superior predictive accuracy and possess a broader spectral radius, while their short‑term memory capacity remains comparable to that of traditional models. Moreover, Roeschies et al. [roeschies2010structure] proposed an evolutionary algorithm that optimizes reservoir‑network architectures for specific problem classes, resulting in substantial gains in predictive performance. Sun et al. [sun2012modeling] introduced a deterministic echo state network architecture featuring a loop reservoir with adjacent feedback connections. This simplified design builds upon the basic loop reservoir structure by incorporating regular adjacent feedback, requiring optimization of only a single parameter. The approach significantly reduces the complexity of ESN implementation while maintaining computational capabilities. These results, together with previous studies, confirm that the structure of a reservoir is pivotal in determining its functionality and overall effectiveness.
In a broader optimization context, surrogate-based and response surface methodologies (RSM) have been extensively studied in engineering design and system optimization. These approaches typically construct an explicit or implicit approximation of a performance landscape over a continuous parameter space, and subsequently perform optimization on the learned surrogate model. Representative examples include [ahmed2024experimental, ahmed2025experimental], where response surfaces are employed to guide design optimization under computational or experimental constraints. While effective for continuous and parameterized systems, such approaches fundamentally differ from our concern: the reservoir structure considered in this work is discrete, graph-based, and non-differentiable. We will focus on topology-level structural modifications guided by algebraic-topological invariants, providing a complementary optimization paradigm that is particularly suitable for discrete dynamical networks.
@@ 2.2 GLMY homology and its applications
GLMY homology represents a significant extension of classical homology theory, specifically developed to analyze directed or asymmetric interactions in digraphs [grigor2012homologies] and hypergraphs [grigor2019homology]. This framework has spawned several theoretical developments, such as homotopy theory for digraphs [grigor2014homotopy], discrete Morse theory on digraphs [lin2021discrete], and path homology theory for multigraphs and quivers [grigor2018path]. Notable computational advances include the persistent path homology for networks proposed by Chowdhury and Mémoli [chowdhury2018persistent], which offers a novel approach to digraph analysis in topological data analysis, and an efficient algorithm for computing one-dimensional persistent path homology of digraph filtrations by Dey et al. [dey2022efficient]. In contrast to traditional simplicial complexes, GLMY homology offers a unique set of dimensional fingerprints. These fingerprints are instrumental in unearthing the intrinsic structural information hidden within networks, providing a more detailed view of network architecture. This theoretical framework has demonstrated its practical utility across a diverse range of applications, such as modeling intricate biological systems [feng2024hypernetwork], analyzing material properties and structures [liu2023neighborhood], analyzing discrete planar vector fields [chen2025analyzing], and understanding complex networks [chowdhury2019path, chowdhury2020path]. Persistent GLMY homology represents a further advancement in this area. By tracking how topological features change and evolve across various filtration processes, it allows us to identify and analyze topological patterns in networks [chowdhury2018persistent]. This capability significantly enhances our understanding of network behavior and evolution.
@@ 3 Preliminary
A digraph is an ordered pair $G=(V,E)$, where $V$ denotes the vertex set and $E\subseteq V\times V$ represents the directed edges. A digraph without loops or multiple edges is called a simple digraph. The GLMY homology (or path homology) provides a topological framework for analyzing simple digraphs [grigor2012homologies]. For a finite nonempty vertex set $V$, an elementary $n$-path is defined as a sequence $e_{i_{0}\ldots i_{n}}$ of $n+1$ vertices, where $n\geq 0$. By convention, the set of elementary (1)(-1)-paths is empty. The boundary operator acts on elementary $n$-paths as:
@@
Figure 2: (a) Example of a one-dimensional persistence diagram. (b) An illustration of the confidence set (red area) of (a).
Persistent homology is a tool from computational topology that captures the topological features of space at various scales [edelsbrunner2008persistent]. It provides a way to quantify the persistence of topological features, capturing the birth and death of topological features in the considered space (such as point clouds, graphs, and digraphs) by constructing a filtration. The persistence diagram (PD) serves as a powerful visualization tool for persistent homology, where each point $(b_{i},d_{i})$ corresponds to a persistence pair representing the birth ($b_{i}$) and death ($d_{i}$) times of a homology generator (Fig. 2). Notice that all points in a PD are located above or on the diagonal. And the value $|d_{i}-b_{i}|$ is called the persistence of the point $(b_{i},d_{i})$, representing the significance of the corresponding topological features. The greater the persistence of a point in the persistence diagram, the more significant the corresponding topological feature is.
A statistical approach for identifying significant points in a persistence diagram is the confidence set [fasy2014confidence], which categorizes these points into two classes based on their persistence. As shown in Fig. 2 (b), the confidence set draws two areas on the persistence diagram: the red area and the white area, which contain points with less persistence and greater persistence, respectively. Points in the class that show greater persistence are called significant points, while the other points are called noise points.
The tool of persistence diagram is first employed to illustrate the persistent homology computed from some point clouds, and this framework can also naturally extend to persistent GLMY homology, where we denote the one-dimensional persistence diagram of $H_{1}$ as $PD_{1}$.